Let x, y ,z be in A.P. and their sum is 15. If 1 , 3 , 9 are added to these three numbers respectively , then they form a G.P. find x , y and z.


1
Expert's answer
2021-03-10T11:19:15-0500

Let three first numbers in an arithmetic progression are "x, x+d, x+2d."

Then "x+x+d+x+2d=15"


"\dfrac{x+d+3}{x+1}=\dfrac{x+2d+9}{x+d+3}"

"3x+3d=15=>d=5-x"

"(x+5-x+3)^2=(x+1)(x+2(5-x)+9)"

"(x+1)(19-x)=64"

"-x^2+18x+19=64"

"x^2-18x+45=0"

"(x-3)(x-15)=0"

"x_1=3, x_2=15""d_1=2, d_2=-10"

"x=3, y=5, z=7"

or

"x=15, y=5,z=-5"